OpenStudy (anonymous):

What is negative square root 2 plus negative square root 2?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[-\sqrt{2}+(-\sqrt{2})=-\sqrt{2}-\sqrt{2}=-2\sqrt{2}\]

OpenStudy (anonymous):

Thanks you, I was wondering if it was that or \[-1\sqrt{2}\]